Circular No. 8741

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006ed–2006en Several additional apparent supernovae have been discovered on unfil-

tered CCD images: 2006en by T. Puckett and M. Peoples (0.60-m reflector at Ellijay, GA; cf. IAUC 8733), and the rest by N. Joubert, D. R. Madison, R. Mostardi, H. Khandrika, and W. Li (LOSS/KAIT; cf. IAUC 8736).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ed Aug. 18.17 15^{h}03^{m}50.^{s}64 +42^{o}06′48′′.5 17.7 1′′.8 E, 7′′.2 S
2006ee Aug. 18.47 1 59 35.40 +14 00 19.6 17.6 9′′.5 E, 9′′.4 S
2006ef Aug. 18.51 2 04 19.51 − 8 43 42.2 15.6 8′′.3 E, 24′′.9 N
2006eg Aug. 20.51 2 32 53.57 +19 15 33.2 18.3 0′′.6 E, 3′′.2 N
2006eh Aug. 20.52 2 50 11.95 +47 03 20.5 18.2 3′′.2 W, 3′′.6 N
2006ei Aug. 21.48 1 56 37.15 +34 10 37.4 18.5 10′′.3 W, 0′′.1 N
2006ej Aug. 23.39 0 38 59.77 − 9 00 56.6 16.4 6′′.2 W, 4′′.6 S
2006ek Aug. 25.27 22 05 09.23 +26 36 31.4 19.2 3′′.6 E, 5′′.8 N
2006el Aug. 25.28 22 47 38.50 +39 52 27.6 18.2 12′′.9 E, 18′′.2 S
2006em Aug. 25.49 2 25 44.30 +41 56 31.8 17.5 21′′.4 E, 50′′.9 S
2006en Aug. 26.15 23 10 05.06 +30 13 24.0 15.8 10′′.9 E, 3′′.8 S

Additional magnitudes for 2006ed in UGC 9684: Aug. 2.21 UT, [19.0; 5.19, [18.0; 19.19, 17.6. Additional magnitudes for 2006ee in NGC 774: Aug. 10.45, [18.7; 19.44, 17.7. Additional magnitudes for 2006ef in NGC 809: Aug. 1.49, [18.3; 19.41, 15.3. Additional magnitudes for 2006eg: Jan. 20.12, [19.5; July 31.52, [18.5; Aug. 22.50, 18.1. Additional magnitudes for 2006eh in UGC 2308: Feb. 22.12, [19.5; Aug. 2.54, [18.3; 22.51, 18.3. Additional magnitudes for 2006ei in NGC 735: Jan. 30.13, [19.5; July 13.49, [18.2; Aug. 4.49, [18.0; 22.51, 18.2. Additional magnitudes for 2006ej in NGC 191A: Aug. 14.40, [19.3; 24.38, 14.7. NGC 191A is the southeast component of the interacting galaxy pair Arp 127 (the other component being NGC 191). Additional magnitudes for 2006ek in MCG +04-52-3: Aug. 17.39 UT, [19.5; 26.30, 19.0. Additional magnitudes for 2006el in UGC 12188: Aug. 18.32, [19.4; 26.30, 18.0. SN 2006dn also appeared in UGC 12188. Additional magnitudes for SN 2006em in NGC 911: Aug. 19.53, [19.2; 26.44, 17.4. Additional magnitudes for 2006en in MCG +05-54-41: 2005 Dec. 15, [19.3; 2006 Aug. 27.19, 15.2 (note that the dates given here for 2006en correct those given on CBET 606). Spectra by R. J. Foley et al. on Aug. 24 UT (cf. CBET 604) yield the following supernova types: 2006ed, II; 2006ee, II; 2006ef, Ia; 2006eg, Ib/c (nebular phase); 2006eh, IIn, 2006ei, Ic; 2006ej, Ia.

2006 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green